# Geometry and Optimization of Shallow Polynomial Networks

### v1

@@ 2.1 The functional space
${\rm Crit}(\tau_{r})=\left\{\mathcal{W}\colon{\rm rk}(d\tau(\mathcal{W}))<\dim(\mathcal{F}_{r})\right\}\subset\mathbb{R}^{r}\times\mathbb{R}^{r\times n}.$
@@
When $r\geq r_{\rm thick}(d,n)$ the critical parameter set ${\rm Crit}(\tau_{r})$ can be seen as a low-dimensional subset of “bad” parameters where all critical points except global minima are confined. Importantly, this set does not depend on the particular functions $L$ or $\ell$ but only on the architecture of the model. This provides a criterion for whether a critical point $\mathcal{W}^{*}$ reached by a local optimization method is a global minimum: if $\mathcal{W}^{*}\not\in{\rm Crit}(\tau_{r})$, it is necessarily a global minimum; if $\mathcal{W}^{*}\in{\rm Crit}(\tau_{r})$, then then it is a not a global minimum unless $\ell$ admits a global minimum on ${\rm Crit}(\tau_{r})$ (which would not occur for generic data).
@@
The second claim of Proposition 6 is similar to results that have previously appeared (see, e.g., [21, Theorem 2.4]), although these works do not mention the critical locus of the parameterization. To the best of our knowledge, the following result is new.
@@
If ${\rm rk}(d\tau_{r}(\mathcal{W}))<\dim({\rm Sym}^{d}(\mathbb{R}^{n}))$ then $V\neq{\rm Sym}^{d}(\mathbb{R}^{n})$ and there exists $P\in{\rm Sym}^{d}(\mathbb{R}^{n})\setminus\{0\}$ such that $P\in V^{\perp}$, where $V^{\perp}$ is the orthogonal space to $V$ with respect to the Frobenius inner product. It is now sufficient to note that $\langle P,{\rm sym}(x^{\otimes(d-1)}\otimes u)\rangle_{F}=\nabla_{x}\langle P,x^{\otimes d}\rangle\cdot u$. Finally, if $w_{i}=0$ whenever $\alpha_{i}=0$, then $V$ is the image of $d\tau_{r}(\mathcal{W})$, so the existence of $P$ means that ${\rm rk}(d\tau_{r}(\mathcal{W}))<\dim({\rm Sym}^{d}(\mathbb{R}^{n}))$. ∎
@@ 2.2 Optimization landscapes
Most works on the loss landscape of neural networks aim to show the absence of bad local minima in the loss landscape [40, 18, 27]. One exception is [42], which discusses conditions on the width of a network that guarantee that the loss has no spurious valleys. In particular, this work shows the following useful result, here rephrased and adapted to our setting.
Theorem 9 ([42], Theorem 8).
@@
It is important to remark that there is no general implication between the absence of bad local minima and the absence of spurious valleys (see Figure 1). Indeed, a function might have a bad local minimum without a spurious valley, if the minimum is non-strict (Figure 1, (c)). Conversely, and perhaps less obviously, a function can have a spurious valley without a bad local minimum, if the minimum of the valley is at infinity (Figure 1, (d)). [^1] Lemma 2 in [42] makes a small mistake and claims that “presence of spurious valleys implies existence of spurious minima.” This is true only if all minima are finite. We also introduce third notion that is related to the absence of spurious valleys:
  • We say that a point $x\in\mathbb{R}^{N}$ is escapable for a continuous $f:\mathbb{R}^{N}\rightarrow\mathbb{R}$ if there exists a path $\gamma:[0,1]\rightarrow M$ such that $\gamma(0)=x$ along which $f$ does not increase (i.e., $f(\gamma(t_{1}))\geq f(\gamma(t_{2}))$ for all $t_{1}<t_{2}$) and such that $f(\gamma(1))<f(x)$. Note that if $f$ is smooth, any point that does not belong to ${\rm Crit}(f)$ is escapable.
A function that has no spurious valleys is such that all points in its domain are escapable. However, there exist functions that have spurious valleys but for which all points are escapable. An example is the function shown in Figure 1 (d).
@@
We refine some results from [42] that relate the optimization landscape to network width. We start by discussing a negative result: when the network architecture is not large enough, bad local minima may exist. According to [42, Corollary 14], if $d$ is even and $L$ is the quadratic loss, then $L$ may have spurious valleys when $r\leq\frac{1}{2}r_{\rm fill}(d,n-1)$ if $d>2$ and when $r\leq n$ if $d=2$. We provide a more general result that considers arbitrary loss functions and applies to even and odd $d$.
@@
If $cl(\mathcal{F}_{r})\neq{\rm Sym}^{d}(\mathbb{R}^{n})$, then $cl(\mathcal{F}_{r})$ is not a convex subset of ${\rm Sym}^{d}(\mathbb{R}^{n})$. Indeed, for any $T\in{\rm Sym}^{d}(\mathbb{R}^{n})\setminus cl(\mathcal{F}_{r})$ that decomposes as $T=\sum_{i=1}^{s}\alpha_{i}w_{i}^{\otimes d}$ (with $s>r$), we have that $T=\frac{1}{s}\sum_{i=1}^{s}s\alpha_{i}w_{i}^{\otimes d}$ is a convex combination of elements in $\mathcal{F}_{1}\subset\mathcal{F}_{r}$. The fact that $cl(\mathcal{F}_{r})$ is not convex easily implies that we can find a convex function $\ell:{\rm Sym}^{d}(\mathbb{R}^{n})\rightarrow\mathbb{R}^{n}$ whose restriction to $\mathcal{F}_{r}$ has (arbitrarily bad) minima and spurious valleys (see [30, Proposition 6]). ∎
Note that $\frac{1}{2}r_{\rm fill}(d,n-1)\leq r_{\rm thick}(d,n-1)<r_{\rm thick}(d,n)$ so compared to [42] we show that bad minima may exist for larger networks. On the positive side — when the network is wide the landscape is favorable — the following two results improve Theorem 9 in two special cases.
@@
Assume $n=2$ and $r\geq\lceil\frac{d+1}{2}\rceil$ (recall $r_{\rm thick}(d,2)=\lceil\frac{d+1}{2}\rceil$). Then any function of the form $L=\tau_{r}\circ\ell$ where $\ell:{\rm Sym}^{d}(\mathbb{R}^{2})$ is smooth and convex is such that all points are escapable.
@@
Assume $d=2$ and $r\geq n$ (recall $r_{\rm fill}(2,n)=r_{\rm thick}(2,n)=n$). Then any function of the form $L=\tau_{r}\circ\ell$ where $\ell:{\rm Sym}^{2}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ is smooth and convex has no spurious valleys.
@@
For $n=2$, it is known that $cl(\mathcal{F}_{r})\neq{\rm Sym}^{d}(\mathbb{R}^{n})$ holds if and only if $r<d$ (this is because all ranks between $\lceil\frac{d+1}{2}\rceil$ and $d$ are “typical” [9, Theorem 2.4]). Combining Theorem 11 and Proposition 10, we deduce that for $n=2$ and $\lceil\frac{d+1}{2}\rceil\leq r<d$ there exist convex functions $\ell:{\rm Sym}^{d}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ with spurious valleys, however even for such functions all (critical) points are escapable. The landscape in these cases is analogous to the one shown in Figure 1, (d). On the other hand, not all convex functions $\ell:{\rm Sym}^{d}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ arise as standard loss functions such as the quadratic loss considered in the next sections. It would be interesting to find concrete examples of functions with this kind of landscape (see also the related Example 19 below).
@@
Before arguing these two points we observe that they imply the claim. Indeed, let $\mathcal{W}_{\min}$ be a global minimum of $L$ and let $H\subset\mathbb{R}^{r}\times\mathbb{R}^{r\times n}$ be a connected component of a sublevel set of $L$ such that $\lim\inf_{\mathcal{W}\in H}L(\mathcal{W})>L(\mathcal{W}_{\min})$ (so $H$ is a spurious valley). Let $V$ be a neighborhood of $\mathcal{W}_{\min}$ such that $L(\mathcal{W}^{\prime})<\lim\inf_{\mathcal{W}\in H}L(\mathcal{W})$ for all $\mathcal{W}^{\prime}\in V$. The image of $V$ and $H$ under $\tau_{r}$ are full measure sets in ${\rm Sym}^{2}(\mathbb{R}^{n})$ (by the strong version of Sard’s Theorem). [^2] We can assume without loss of generality that $H$ is open, by replacing a closed sublevel set $L^{-1}([-\infty,t])$ with an open sublevel set $L^{-1}([-\infty,t+\epsilon))$ for $\epsilon$ sufficiently small. Using the second point above, there exists $S_{1}\in\tau_{r}(H)$ and $S_{2}\in\tau_{r}(V)$ such that the path $\gamma(t)=(1-t)S_{1}+tS_{2}$ does not contain matrices with repeated eigenvalues. By 1, this path can be lifted to a path $\mathcal{W}(t)$ such that $\mathcal{W}(0)\in H$, and $\tau_{r}(\mathcal{W}(t))=\gamma(t)$. Since $\ell$ is convex, $\ell(\gamma(t))$ is either monotone decreasing between 00 and $1$ (because $\ell(S_{1})>\ell(S_{2})$ by assumption), or it achieves a global minimum at $t_{0}\in(0,1)$, where the value is less than $\ell(S_{2})$. Either way, the lifted path $\mathcal{W}(t)$ for $t\in[0,1]$ or $t\in[0,t_{0}]$ is a descent path from a point on $H$ to a point where $L$ is less than $\lim\inf_{\mathcal{W}\in H}L(\mathcal{W})$, which is a contradiction.
The two points above follow from known properties of symmetric matrices and their eigenvalues. The first point can be argued by noting that when eigenvalues are distinct, both eigenvalues and eigenvectors are described locally by continuous functions (by the implicit function theorem, see also [1, Section 7]). The second point follows from the fact that the set of (real) matrices with repeated eigenvalues is an algebraic set of codimension two in ${\rm Sym}^{2}(\mathbb{R}^{n})$ [41, Section 7.5]. In particular, a generic line will not meet this set. ∎
@@
We now observe that the loss landscape can have bad local minima for networks with arbitrarily large width $r$. Since by Theorem 9, when $r\geq\dim{\rm Sym}^{d}(\mathbb{R}^{n})$ the landscape has no spurious valleys, we deduce that any bad local minimum must be escapable. The landscape in this case resembles the one shown in Figure 1, (c). A similar example was given for the case $d=2$ (quadratic activations) in [28].
@@
Let $d=2k$ and choose $r\geq r_{\rm fill}$ be arbitrarily large. Consider a function $L$ of the form
@@
Compared to [28], we also show that these bad minima are relevant for gradient-based optimization, since they have a basin of attraction of positive measure. We will use the following property of gradient flow for polynomial networks, which follows from the homogeneity of the activation function. An analogous result has been shown for ReLU networks [19].
@@
for all $t$ and $i$, from the chain rule. Here $\nabla\ell$ is short for $\nabla\ell\left(\tau_{r}(\alpha(t),W(t))\right)$ in tensorized form (so it lies in $\textup{Sym}^{d}(\mathbb{R}^{n})$). Also, $\cdot$ stands for contraction. Eliminating $\nabla\ell$,
@@
for all $t$ and $i$. The result is immediate by anti-differentiating (8) ∎
@@
Assume $d$ even and $r\geq r_{\rm fill}$ arbitrarily large. Let $\ell:{\rm Sym}^{d}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ be a smooth convex function with an isolated global minimum at $T\in{\rm Sym}^{d}(\mathbb{R}^{n})$ such that $g(X)=\langle T,X^{\otimes d}\rangle_{F}$ is a negative polynomial. In particular, $\ell$ could be of the form
@@
The target $T$ lies in the functional space $\mathcal{F}_{r}={\rm Sym}^{d}(\mathbb{R}^{n})$ but is not dynamically reachable. Indeed, it follows from Proposition 16 that $\alpha_{i}>0$ holds throughout the trajectory initialized at $\mathcal{W}_{0}$. Every polynomial in the trajectory is positive, so $\ell$ never reaches zero. ∎
@@
Corollary 17 implies that for polynomial networks of even degree, gradient flow can fail to reach global optima for certain initializations, regardless of the network’s width. However, with random initializations where the top coefficient $\alpha_{i}$ has an equal likelihood of being positive or negative, the condition $\alpha_{i}^{2}>\frac{1}{d}|w_{i}|^{2}$ becomes increasingly unlikely as the width $r$ increases. Therefore, although the optimization landscape is never trivial for wide networks, favorable initializations become more probable.
We conclude this section with an example in which the the global minimum of $L=\ell\circ\tau_{r}$ is at infinity in parameter space, even though $\ell:{\rm Sym}^{d}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ has compact sublevel sets. This phenomenon is due to the fact that the set of low-rank tensors is not topologically closed for $d\geq 3$ (see [14, Example 6.6]).
@@
[…]

### v2

@@ 2.1 The function space
${\rm Crit}(\tau_{r})=\left\{\mathcal{W}\colon{\rm rk}(d\tau_{{r}}(\mathcal{W}))<\dim(\mathcal{F}_{r})\right\}\subset\mathbb{R}^{r}\times\mathbb{R}^{r\times n}.$
@@
When $r\geq r_{\rm thick}(d,n)$ the critical parameter set ${\rm Crit}(\tau_{r})$ can be seen as a low-dimensional subset of “bad” parameters where all critical points except global minima are confined. Importantly, this set does not depend on the particular functions $L$ or $\ell$ but only on the architecture of the model. This provides a criterion for whether a critical point $\mathcal{W}^{*}$ reached by a local optimization method is a global minimum: if $\mathcal{W}^{*}\not\in{\rm Crit}(\tau_{r})$, it is necessarily a global minimum; if $\mathcal{W}^{*}\in{\rm Crit}(\tau_{r})$, then it is a not a global minimum unless $\ell$ admits a global minimum on ${\rm Crit}(\tau_{r})$ (which would not occur for generic data).
@@
The second claim of Proposition 6 is similar to results that have previously appeared (see, e.g., [gamarnik2020, Theorem 2.4]), although these works do not mention the critical locus of the parameterization. To the best of our knowledge, the following result is new.
@@
If ${\rm rk}(d\tau_{r}(\mathcal{W}))<\dim({\rm Sym}^{d}(\mathbb{R}^{n}))$ then $V\neq{\rm Sym}^{d}(\mathbb{R}^{n})$ and there exists $P\in{\rm Sym}^{d}(\mathbb{R}^{n})\setminus\{0\}$ such that $P\in V^{\perp}$, where $V^{\perp}$ is the orthogonal space to $V$ with respect to the Frobenius inner product. It is now sufficient to note that $\langle P,{\rm sym}({X}^{\otimes(d-1)}\otimes u)\rangle_{F}=\nabla_{{X}}\langle P,{X}^{\otimes d}\rangle_{{F}}\cdot u$. Finally, if $w_{i}=0$ whenever $\alpha_{i}=0$, then $V$ is the image of $d\tau_{r}(\mathcal{W})$, so the existence of $P$ means that ${\rm rk}(d\tau_{r}(\mathcal{W}))<\dim({\rm Sym}^{d}(\mathbb{R}^{n}))$. ∎
@@ 2.2 Optimization landscapes
The absence of spurious valleys ensures that from any point in the parameter space there exists a path to a global minimum along which the loss is non-increasing, a property that can be beneficial for gradient-based or other local search methods. Most works on the loss landscape of neural networks aim to show the absence of bad local minima in the loss landscape [soltanolkotabi2018, du2018d, kawaguchi2016]. One exception is [venturi2018spurious], which discusses conditions on the width of a network that guarantee that the loss has no spurious valleys. In particular, this work shows the following useful result, here rephrased and adapted to our setting.
Theorem 9 ([venturi2018spurious], Theorem 8).
@@
It is important to remark that there is no general implication between the absence of bad local minima and the absence of spurious valleys (see Figure 1). Indeed, a function might have a bad local minimum without a spurious valley, if the minimum is non-strict (Figure 1, (c)). Conversely, and perhaps less obviously, a function can have a spurious valley without a bad local minimum, if the minimum of the valley is at infinity (Figure 1, (d)). [^1] Lemma 2 in [venturi2018spurious] makes a small mistake and claims that “presence of spurious valleys implies existence of spurious minima.” This is true only if all minima are finite. We also introduce a third notion that is related to the absence of spurious valleys:
  • We say that a point $x\in\mathbb{R}^{N}$ is escapable for a continuous $f:\mathbb{R}^{N}\rightarrow\mathbb{R}$ if there exists a path $\gamma:[0,1]\rightarrow{\mathbb{R}^{N}}$ such that $\gamma(0)=x$ along which $f$ does not increase (i.e., $f(\gamma(t_{1}))\geq f(\gamma(t_{2}))$ for all $t_{1}<t_{2}$) and such that $f(\gamma(1))<f(x)$. Note that if $f$ is smooth, any point that does not belong to ${\rm Crit}(f)$ is escapable.
A function that has no spurious valleys is such that all points in its domain are escapable or global minima. However, there exist functions that have spurious valleys but for which all points are escapable or global minima. An example is the function shown in Figure 1 (d).
@@
We refine some results from [venturi2018spurious] that relate the optimization landscape to network width. We start by discussing a negative result: when the network architecture is not large enough, bad local minima may exist. According to [venturi2018spurious, Corollary 14], if $d$ is even and $L$ is the quadratic loss, then $L$ may have spurious valleys when $r\leq\frac{1}{2}r_{\rm fill}(d,n-1)$ if $d>2$ and when $r\leq n$ if $d=2$. We provide a more general result that considers arbitrary loss functions and applies to even and odd $d$.
@@
If $cl(\mathcal{F}_{r})\neq{\rm Sym}^{d}(\mathbb{R}^{n})$, then $cl(\mathcal{F}_{r})$ is not a convex subset of ${\rm Sym}^{d}(\mathbb{R}^{n})$. Indeed, for any $T\in{\rm Sym}^{d}(\mathbb{R}^{n})\setminus cl(\mathcal{F}_{r})$ that decomposes as $T=\sum_{i=1}^{s}\alpha_{i}w_{i}^{\otimes d}$ (with $s>r$), we have that $T=\frac{1}{s}\sum_{i=1}^{s}s\alpha_{i}w_{i}^{\otimes d}$ is a convex combination of elements in $\mathcal{F}_{1}\subset\mathcal{F}_{r}$. The fact that $cl(\mathcal{F}_{r})$ is not convex easily implies that we can find a convex function $\ell:{\rm Sym}^{d}(\mathbb{R}^{n})\rightarrow\mathbb{R}^{n}$ whose restriction to $\mathcal{F}_{r}$ has (arbitrarily bad) minima and spurious valleys (see [kileel2019expressive, Proposition 6]). ∎
Note that $\frac{1}{2}r_{\rm fill}(d,n-1)\leq r_{\rm thick}(d,n-1)<r_{\rm thick}(d,n)$ so compared to [venturi2018spurious] we show that bad minima may exist for larger networks. On the positive side — when the network is wide the landscape is favorable — the following two results improve Theorem 9 in two special cases.
@@
Assume $n=2$ and $r\geq\lceil\frac{d+1}{2}\rceil$ (recall $r_{\rm thick}(d,2)=\lceil\frac{d+1}{2}\rceil$). Then any function of the form $L={\ell\circ\tau_{r}}$ where $\ell:{\rm Sym}^{d}(\mathbb{R}^{2})$ is smooth and convex is such that all points that are not global minima are escapable.
@@
Assume $d=2$ and $r\geq n$ (recall $r_{\rm fill}(2,n)=r_{\rm thick}(2,n)=n$). Then any function of the form $L={\ell\circ\tau_{r}}$ where $\ell:{\rm Sym}^{2}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ is smooth and convex has no spurious valleys.
@@
For $n=2$, it is known that $cl(\mathcal{F}_{r})\neq{\rm Sym}^{d}(\mathbb{R}^{n})$ holds if and only if $r<d$ (this is because all ranks between $\lceil\frac{d+1}{2}\rceil$ and $d$ are “typical” [blekherman2012, Theorem 2.4]). Combining Theorem 11 and Proposition 10, we deduce that for $n=2$ and $\lceil\frac{d+1}{2}\rceil\leq r<d$ there exist convex functions $\ell:{\rm Sym}^{d}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ with spurious valleys, however even for such functions all (critical) points are escapable. The landscape in these cases is analogous to the one shown in Figure 1, (d). On the other hand, not all convex functions $\ell:{\rm Sym}^{d}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ arise as standard loss functions such as the quadratic loss considered in the next sections. It would be interesting to find concrete examples of functions with this kind of landscape (see also the related Example 19 below).
@@
Before arguing these two points we observe that they imply the claim. Indeed, let $\mathcal{W}_{\min}$ be a global minimum of $L$ and let $H\subset\mathbb{R}^{r}\times\mathbb{R}^{r\times n}$ be a connected component of a sublevel set of $L$ such that $\lim\inf_{\mathcal{W}\in H}L(\mathcal{W})>L(\mathcal{W}_{\min})$ (so $H$ is a spurious valley). Let $V$ be a neighborhood of $\mathcal{W}_{\min}$ such that $L(\mathcal{W}^{\prime})<\lim\inf_{\mathcal{W}\in H}L(\mathcal{W})$ for all $\mathcal{W}^{\prime}\in V$. The image of $V$ and $H$ under $\tau_{r}$ are full measure sets in ${\rm Sym}^{2}(\mathbb{R}^{n})$ (by the strong version of Sard’s Theorem). [^2] We can assume without loss of generality that $H$ is open, by replacing a closed sublevel set $L^{-1}([-\infty,t])$ with an open sublevel set $L^{-1}([-\infty,t+\epsilon))$ for $\epsilon$ sufficiently small. Using the second point above, there exists $S_{1}\in\tau_{r}(H)$ and $S_{2}\in\tau_{r}(V)$ such that the path $\gamma(t)=(1-t)S_{1}+tS_{2}$ does not contain matrices with repeated eigenvalues. By 1, this path can be lifted to a path $\mathcal{W}(t)$ such that $\mathcal{W}(0)\in H$, and $\tau_{r}(\mathcal{W}(t))=\gamma(t)$. Since $\ell$ is convex, $\ell(\gamma(t))$ is either monotone decreasing between 0 and $1$ (because $\ell(S_{1})>\ell(S_{2})$ by assumption), or it achieves a global minimum at $t_{0}\in(0,1)$, where the value is less than $\ell(S_{2})$. Either way, the lifted path $\mathcal{W}(t)$ for $t\in[0,1]$ or $t\in[0,t_{0}]$ is a descent path from a point on $H$ to a point where $L$ is less than $\lim\inf_{\mathcal{W}\in H}L(\mathcal{W})$, which is a contradiction.
The two points above follow from known properties of symmetric matrices and their eigenvalues. The first point can be argued by noting that when eigenvalues are distinct, both eigenvalues and eigenvectors are described locally by continuous functions (by the implicit function theorem, see also [alekseevsky1998, Section 7]). The second point follows from the fact that the set of (real) matrices with repeated eigenvalues is an algebraic set of codimension two in ${\rm Sym}^{2}(\mathbb{R}^{n})$ [sturmfels, Section 7.5]. In particular, a generic line will not meet this set. ∎
@@
We now observe that the loss landscape can have bad local minima for networks with arbitrarily large width $r$. Since by Theorem 9, when $r\geq\dim{\rm Sym}^{d}(\mathbb{R}^{n})$ the landscape has no spurious valleys, we deduce that any bad local minimum must be escapable. The landscape in this case resembles the one shown in Figure 1, (c). A similar example was given for the case $d=2$ (quadratic activations) in [kazemipour2019no].
@@
Let $d=2k$ and choose $r\geq r_{\rm fill}$ to be arbitrarily large. Consider a function $L$ of the form
@@
Compared to [kazemipour2019no], we also show that these bad minima are relevant for gradient-based optimization, since they have a basin of attraction of positive measure. We will use the following property of gradient flow for polynomial networks, which follows from the homogeneity of the activation function. An analogous result has been shown for ReLU networks [du2018algorithmic].
@@
for all $t$ and $i$, from the chain rule (see equation (6)). Here $\nabla\ell$ is short for $\nabla\ell\left(\tau_{r}(\alpha(t),W(t))\right)$ in tensorized form (so it can be viewed as a tensor in $\textup{Sym}^{d}(\mathbb{R}^{n})$). Also, $\cdot$ stands for contraction. Eliminating $\nabla\ell$,
@@
for all $t$ and $i$. The result is immediate by anti-differentiating (8). ∎
@@
Assume $d$ is even and $r\geq r_{\rm fill}$ is arbitrarily large. Let $\ell:{\rm Sym}^{d}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ be a smooth convex function with an isolated global minimum at $T\in{\rm Sym}^{d}(\mathbb{R}^{n})$ such that $g(X)=\langle T,X^{\otimes d}\rangle_{F}$ is a negative polynomial. In particular, $\ell$ could be of the form
@@
The target $T$ lies in the function space $\mathcal{F}_{r}={\rm Sym}^{d}(\mathbb{R}^{n})$ but is not dynamically reachable. Indeed, it follows from Proposition 16 that $\alpha_{i}>0$ holds throughout the trajectory initialized at $\mathcal{W}_{0}$. Every polynomial in the trajectory is positive, so $\ell$ never reaches zero. ∎
@@
Corollary 17 implies that for polynomial networks of even degree, gradient flow can fail to reach global optima for certain initializations, regardless of the network’s width. However, with random initializations where the top coefficient $\alpha_{i}$ has an equal likelihood of being positive or negative, the condition $\alpha_{i}^{2}>\frac{1}{d}\|w_{i}\|^{2}$ for all ii—which imposes that all dyanmically reachable polynomials are positive—becomes increasingly unlikely at initialization as the width $r$ increases. Therefore, although the optimization landscape is never trivial for wide networks, favorable initializations become more probable.
We conclude this section with an example in which the the global minimum of $L=\ell\circ\tau_{r}$ is at infinity in parameter space, even though $\ell:{\rm Sym}^{d}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ has compact sublevel sets. This phenomenon is due to the fact that the set of low-rank tensors is not topologically closed for $d\geq 3$ (see [comon2008, Example 6.6]).
@@
Consider the teacher model $g(x)=x_{1}^{d-1}x_{2}$ and let $L(\mathcal{W})=\|f_{\mathcal{W}}(x)-g(x)\|^{2}$ for some norm $\|\cdot\|$ on ${\rm Sym}^{d}(\mathbb{R}^{n})$. It is possible to show that $f_{\mathcal{W}}\neq g$ holds for all parameters $\mathcal{W}\in\mathbb{R}^{2}\times\mathbb{R}^{2\times n}$. Indeed, $f_{\mathcal{W}}$ corresponds to a tensor of rank at most $2$, while $g(x)$ corresponds to a tensor of rank exactly $d$ (see Proposition 5.6 in [comon2008]). It follows that $L(\mathcal{W})$ is always strictly positive. However, consider the following sequence of parameters $\mathcal{W}(\tau)$ for $\tau\rightarrow\infty$: